OpenStudy (anonymous):

how would you solve x/x^2-4+2/x-2=x-3/x^2-x-6?

OpenStudy (mertsj):

\[\frac{x}{x^2-4}+\frac{2}{x-2}=\frac{x-3}{x^2-x-6}\]

OpenStudy (mertsj):

1. Factor each denominator 2. Reduce the fraction on the right 3. Multiply each term by the common denominator 4. Solve the resulting equation.